# Consequences of a Two-Time Relativistic Bohmian Model

### v14

@@ IV.3 Standard Zitterbewegung
Interpretating the oscillatory motion in $\tau$ as occurring in t, the standard Zitterbewegung is re-obtained. In the standard view, $R$ is considered time independent, so the second side of eq. (10) is just $-\frac{H^{2}}{\hbar{}^{2}}$, with $H=const$. In the new theory this means demanding: $\frac{1}{R}\frac{\partial^{2}R}{\partial t^{2}}-2\,\frac{i}{\hbar}\,\frac{H}{R}\frac{\partial R}{\partial t}=0$, that provides $R\propto\,e^{\frac{2i}{\hbar}H\,t}$ and hence $V_{Q}\propto e^{-\frac{4i}{\hbar}H\,t}$. Then, using eqs. (22) for $V_{Q}$ and (24) (both as for t-time), after having conveniently determined the arbitrary constants, one can get: $v_{i}\propto c\,e^{-\frac{2i}{\hbar}H\,t}$, as found in the Heisenberg picture.
V Working account for the double-slit experiment
@@
The standard wave packet description, compliant the de Broglie relation, is still empirically justified on the basis of the physical indeterminations of the sent particle. At the initial instant, we assume for simplicity that the particle has the minimum positional uncertainty previously found, which we call $\Delta s_{0}$. During motion, as long as the energy of the particle does not change, the growth of the positional uncertainty, obtainable by the Liouville’s theorem, continues to be the standard one: $\Delta s=\sqrt{\Delta s_{0}^{2}+\frac{\Delta p_{s0}^{2}}{m^{2}}\,t^{2}}$. At a fixed instant $\bar{t}$, $\Delta s$ represents the radius of a sphere-like figure within which the particle, due to the omnidirectional $\tau$-motions, is spread; in other words, at instant $\bar{t}$ it is simultaneously present in every point of this volume. We highlight that:
@@
    This simultaneity is not in conflict with Special Relativity, because the generic $\tau$-motion itself, as long as the energy of the particle does not change, does not constitute an event: $\bar{t}$ is fixed in any reference system.
@@
    In addition to interacting like particles, self-particles interfere like waves, exactly as predicted by the standard interpretation, according to the Feynman interpretation [67]. The model simply specifies that the different histories are realized in time $\tau$. This description is in conflict neither with the interactions just discussed nor with the corpuscular description given by equations (20) and (24). Due to the constitutional limitations of the latter, the effective trajectory followed by the particle is everywhere unknowable and the wave description is necessary. When the particle reaches the wall, diffraction is observed if the sphere-like volume includes the two slits (Fig. 1).
@@
Note that zig-zag trajectories which violate the conservation of $t$-momentum and yet admitted as legitimate Feynman paths, can now be properly physical, being able to be obtained by composition of $\vec{p}_{c}$ with one or more $\vec{p}_{is}$.
@@ V.1 Collapses
Areas where the self-interference are destructive constitute ”shadow zones” of the sphere-like volume, where from a corpuscular point of view, the Quantum Potential pushes the self-particles elsewhere. Dually, there will also be areas where the interference is constructive.
The apparent observational ”collapse” is an extreme case of destructive interference that occurs after a energy variation of the particle for collision (more generally, interaction) with an external particle (or field): it happens that the self-wavefunctions in the impact area change their phase; then, their interference with the self-wavefunctions located elsewhere cancels - instantaneously, with respect to $t$-time - the latter. From a corpuscular point of view, the Quantum Potential no longer allows the particle to $\tau$-move away from the proximity of the impact.
So, a non-zero energetic interaction with an external particle or field constitutes an observable (but not necessarily observed) event, capable of dramatically restricting $\Delta s$, its minimum value being of the Compton length order. From a macroscopic view, the intrinsic movements stop. Observation itself has no special role: this happens regardless of whether the external particle is from a detector or residual air.
@@
This picture also indicates another fundamental reason that generally prevents a macroscopic body from spread out in space through intrinsic motions: compared to a particle, it is vastly more likely that it interacts energetically with other matter or fields. Indeed, even a very small variation in the phase of the wave is capable of making it collapse in the way described, canceling out significant quantum uncertainties.
For what has been said, the paradox that the particle, being $t$-simultaneously in every point of a certain macroscopic volume, violates the conservation of mass and of number of particles is only apparent: a particle search experiment will always detect only one particle. The total purely quantum energy distributed in the sphere-like volume in a $\Delta\tau$ interval is indeed infinite, but it is beyond any kind of experimentation. Really it is virtual, in the same sense used in standard view. Indeed, the existence of the hidden time $\tau$ is capable of explaining why the increasingly larger energies predicted by entering the infinitesimal are real (as experiments excellently confirm) but never directly observed.
@@ VI Discussion
In Feynman’s (however standard) formulation of Quantum Mechanics, when a particle moves from one point to another, it is as if it followed at once all possible paths between the two points. In this article we have begun to explore a model in which, by abandoning the ”as if”, the particle really does so. For this to be possible without contradiction, it is necessary to add a new time parameter, totally independent of the commonly experienced one [^2] After all, time could logically be defined as the parameter that allows a particle to be at different points.. Called it $\tau$, the wavefunction of the particle is represented by eq. (18). In it, the interpretation of S does not change, so all equations referring to time t, as they use partial derivatives, consist just of a relativistic generalization of the classical Bohmian ones. Conversely, by assuming for the generic $\tau$-motion an appropriate continuity equation, new laws of motion, which must replace that one of Bohm’s original theory, derive. The $\tau$-motions are governed by the Quantum Potential and spread out the particle in a region of space that depends on the conditions of the physical surrounding. For a free particle, a more general Zitterbewegung motion has been found, having minimum amplitude of the order of the Compton length. In presence of a $V\neq 0$, the Quantum Potential is modified through eqs. (11) and (12) and the spreading through eq. (24); in the case of an atomic electron, for example, the $\tau$-motions justify the instantaneous dispersion of the electron in the orbit (we intend to show this in a subsequent publication).
In the model, the quantum uncertainties and the uncertainty principle itself have a deterministic explanation: they are due to the $\tau$-movements. For the latter, the Special Relativity does not hold, but since they are judged as instantaneous with respect to t-time, measured in any reference system, it is impossible to detect violations of the Lorentz covariance.
@@
The imposition of eq. (22), necessary to deduce the general equation (24), might seem a weakness of the theory. Anyway, it is legitimate that even the generic movement in $\tau$ possesses own equation of continuity, like that one in t. Its peculiarity, to describe a regressive wave, reminds the motion of an antiparticle.
@@
A question concerns the determination of the parameter f. By assuming f = 1/2, one gets, in any direction, a total purely quantum energy $m\gamma_{c}c^{2}$, which could represent a virtual antiparticle. But, naturally, the determination of f should derive by rigorous considerations and it is not excluded that it may depend on the type of particle and/or its properties here not considered.
@@
The evanescence of the uncertainty principle in direction of motion for $v_{c}\rightarrow c$, is recognizable just by observing the general eq. (49): there, a Lorentz factor - equal to 1 only in direction tending to be perpendicular to the rectilinear classic motion - directly reduces the influence of $\hbar$. Then, the spreading region of the particle should be contracted in direction of motion. Believing that the high energy available in today’s accelerators should have already detected deviations based on eq. (52), would be quite hasty. It is necessary to consider not only that high speeds make localization measurements hard but above all that simultaneous measurements of complementary observables are prohibited by the same principle of causality in the Special Relativity. These and several other problems make it very difficult to experimentally verify the uncertainty principle [69, 68].
@@
Compared to spacetime structures planned by the current most accredited theories, introduction of a second time variable is a more radical but also mathematically simpler criterion which, neither logically nor physically, differ in essence from the introduction of an extra spatial dimension. In addition to the explanation of the Zitterbewegung, it predicts new physical effects, permitting, inside the Feynman’s formulation, to clarify the phase of the paths (redefinable as real physical trajectories, although unobservable) and hopefully to represent the sum over histories as a properly mathematical integral.
@@
Conclusively, in the explored theory the apparently paradoxical aspects of Quantum Mechanics are just the effect of the existence of an extra unobservable time dimension. By traveling in a hidden extra-time dimension, you can deterministically explain why a particle or photon is able to interact instantly with more slits in a diffraction, to run across multiple paths between two points, to admit entanglements with other arbitrarily distant particles… All observed facts.

### v16

@@ IV.3 Standard Zitterbewegung
Interpretating the generic oscillatory motion in $\tau$ as occurring in t, the standard Zitterbewegung is re-obtained. In the standard view, $R$ is considered time independent, so the second side of eq. (10) is just $-\frac{H^{2}}{\hbar{}^{2}}$, with $H=const$. In the new theory this means demanding: $\frac{1}{R}\frac{\partial^{2}R}{\partial t^{2}}-2\,\frac{i}{\hbar}\,\frac{H}{R}\frac{\partial R}{\partial t}=0$, that provides $R\propto\,e^{\frac{2i}{\hbar}H\,t}$ and hence $V_{Q}\propto e^{-\frac{4i}{\hbar}H\,t}$. Then, using eqs. (22) for $V_{Q}$ and (24) (both as for t-time), after having conveniently determined the arbitrary constants, one can get: $v_{i_{s}}\propto c\,e^{-\frac{2i}{\hbar}H\,t}$, as found in the Heisenberg picture.
V Interpretation of the double-slit experiment
@@
The standard wave packet description, compliant the de Broglie relation, is still empirically justified on the basis of the physical indeterminations of the sent particle. At the initial instant, we assume for simplicity that the particle has the minimum positional uncertainty previously found, which we call $\Delta s_{0}$. During motion, as long as the energy of the particle does not change, the growth of the positional uncertainty, obtainable by the Liouville’s theorem, continues to be the standard one: $\Delta s=\sqrt{\Delta s_{0}^{2}+\frac{\Delta p_{s_{0}}^{2}}{m^{2}}\,t^{2}}$. At a fixed instant $\bar{t}$, $\Delta s$ represents the radius of a sphere-like figure within which the particle, due to the omnidirectional $\tau$-motions, is spread; in other words, at instant $\bar{t}$ it is simultaneously present in every point of this volume. We highlight that:
@@
    This simultaneity is not in conflict with Special Relativity, because the generic $\tau$-motion itself, as long as the energy of the particle does not change, does not constitute a $t$-event: $\bar{t}$ is frozen in any reference system.
@@
    In addition to interacting like particles, self-particles interfere like waves, exactly as predicted by the standard interpretation, according to the Feynman description [68]. The model simply specifies that the different histories are actually realized in time $\tau$. This description is in conflict neither with the interactions just discussed nor with the corpuscular description given by equations (20) and (24). Due to the constitutional limitations of the latter, the effective trajectory followed by the particle is everywhere unknowable and the wave description is then necessary. When the particle reaches the wall, diffraction is observed if the sphere-like volume includes the two slits (Fig. 1).
@@
Note that zig-zag trajectories which violate the conservation of $t$-momentum and yet admitted as legitimate Feynman paths, can now be properly physical, being able to be obtained by composition of $\vec{p}_{c}$ with one or more $\vec{p}_{i_{s}}$.
@@ V.1 Collapses
Areas where the self-interferences are destructive constitute ”shadow zones” of the sphere-like volume, where from a corpuscular point of view, the Quantum Potential pushes the self-particles elsewhere. Dually, there will also be areas where the interference is constructive.
The apparent observational ”collapse” is an extreme case of destructive interference that occurs after a energy variation of the particle for collision (more generally, interaction) with an external particle or field: what happens is that the self-wavefunction at the point of impact changes abruptly its phase proportional to $S_{t}$; then, its interference with the self-wavefunctions located elsewhere cancels - instantaneously, with respect to $t$-time - the latter. From a corpuscular point of view, the Quantum Potential no longer allows the particle to $\tau$-move away from the proximity of the impact.
So, a non-zero energetic interaction with an external particle or field constitutes an observable (but not necessarily observed) $t$-event, capable of dramatically restricting $\Delta s$, its minimum value being of the Compton length order. From a macroscopic view, the intrinsic movements stop. Observation itself has no special role: this happens regardless of whether the external particle is from a detector or residual air.
@@
This picture also indicates another fundamental reason that generally prevents a macroscopic body from spread out in space through the intrinsic motions: compared to a particle, it is vastly more likely that it interacts energetically with other matter or fields. Indeed, even a small wave phase variation is capable of making it collapse in the way described, canceling out significant quantum uncertainties.
For what has been said, the paradox that the particle, being $t$-simultaneously in every point of a certain macroscopic volume, violates the conservation of mass and of number of particles is only apparent: a particle search experiment will always detect only one particle. The total purely quantum energy distributed in the sphere-like volume at any $t$-instant is indeed infinite, but it is beyond any kind of experimentation. Actually it is virtual, in the same sense used in standard view. Indeed, the existence of the hidden time $\tau$ is capable of explaining why the increasingly larger energies predicted by entering the infinitesimal are real (as experiments excellently confirm) but never directly observed.
@@ VI Discussion
In Feynman’s (however standard) formulation of Quantum Mechanics, when a particle moves from one point to another, it is as if it followed at once all possible paths between the two points. In this article we have begun to explore a model in which, by abandoning the ”as if”, the particle really does so. For this to be possible without contradiction, it is necessary to add a new time parameter, totally independent of the commonly experienced one. Called it $\tau$, the generic wavefunction of the particle is represented by eq. (18). The equations that are obtained by differentiating the latter with respect to time $t$ consist just of a relativistic generalization of the classical Bohmian ones. Conversely, by assuming for the generic $\tau$-motion an appropriate wave equation, new laws of motion, which must replace that one of Bohm’s original theory, derive. The $\tau$-motions are governed by the Quantum Potential and spread out the particle in a region of space that depends on the conditions of the physical surrounding. For a free particle, a more general Zitterbewegung motion has been found, having minimum amplitude of the order of the Compton length. In presence of a $V\neq 0$, the Quantum Potential is modified through eqs. (11) and (12) and the spreading through eq. (24); in the case of an atomic electron, for example, the $\tau$-motions justify the instantaneous dispersion of the electron in the orbit (we intend to show this in a subsequent publication).
In the model, the quantum uncertainties and the uncertainty principle itself have a deterministic explanation: they are due to the $\tau$-movements. For the latter, the Special Relativity cannot hold, but since they are instantaneous with respect to t-time in any reference system, it is impossible to detect violations of the Lorentz covariance.
@@
A question concerns the determination of the parameter f. For example, by assuming f = 1/2, one gets, in any direction, a total purely quantum energy $m\gamma_{c}c^{2}$, equal to that of the particle. However we find no arguments to suggest whether this is the (or an) appropriate value. It’s clear that the determination of f should derive by rigorous considerations and it is not excluded that it may depend on the type of particle and/or its properties here not considered.
@@
The evanescence of the uncertainty principle in direction of motion for $v_{c}\rightarrow c$, is recognizable just by observing the general eq. (51): there, a Lorentz factor - equal to 1 only in direction tending to be perpendicular to the rectilinear classic motion - directly reduces the influence of $\hbar$. Then, the spreading region of the particle should be contracted in direction of motion. Believing that the high energy available in today’s accelerators should have already detected deviations based on eq. (53), would be quite hasty. It is necessary to consider not only that high speeds make localization measurements hard but above all that simultaneous measurements of complementary observables are prohibited by the same principle of causality in Special Relativity. These and several other problems make it very difficult to experimentally verify the uncertainty principle [69, 70].
@@
Compared to spacetime structures planned by the current most accredited theories, introduction of a second time variable is a more radical but also mathematically simpler criterion which, neither logically nor physically, differ in essence from the introduction of an extra spatial dimension. In addition to the explanation of the Zitterbewegung, it predicts new physical effects, permitting, inside the Feynman’s formulation, to clarify the phase of the paths (redefinable as real physical trajectories, although not always observable) and hopefully to represent the sum over histories as a properly mathematical integral.
@@
In the explored theory, the apparently paradoxical aspects of Quantum Mechanics are just the effect of the existence of an extra unobservable time dimension. By traveling in a hidden time, you can deterministically explain why a particle or photon is able to interact instantly with more slits in a diffraction, to run across multiple paths between two points, to admit entanglements with other arbitrarily distant particles, to emit and reabsorb many species of virtual particles… All established facts.
Appendix
In this section we provide a justification of the peculiar wave equation (22), necessary to deduce the $\tau$-motion law. First of all, taking advantage of the reported generality, it is convenient to refer it to $R^{2}$:
$\frac{\partial R^{2}}{\partial\tau}-v_{i_{s}}{\frac{dR^{2}}{ds}}=0$ (54)
When one refers this motion to an oscillation, one realizes that $R^{2}$ is in antiphase - i.e. shifted by half a period - with respect to the particle moving with speed $v_{i_{s}}$. This ”pursuing” reminds the way the Zitterbewegung is interpreted in the standard interpretation: the particle ”chases” its antiparticle (and vice versa) which, according to the uncertainty principle, is created for lengths $\sim\lambda_{c}$ (see e.g. [71]). However, our model makes no reference to antimatter, admitting only positive values for energy. In fact, we are able to justify this equation by referring exclusively to the meaning of the physical quantities involved. According to our model, the operation of finding the particle in a unit volume (whose probability is $R^{2}$) must specify a $\bar{t}$-event. Then, for an amplitude $\sim\lambda_{c}$, $\bar{t}$ has an uncertainty of about half a period of the intrinsic oscillation of the particle, as we have shown. And this effectively means that the particle must perform half an oscillation before the measurement is achievable, ending up in antiphase with $R^{2}$.